OpenStudy (bigbosssaint21):

complete the square P(n) = -250n^2 + 3,250n - 9,000

OpenStudy (anonymous):

\[-250(n^2-13n+36)\\ -250\left(n^2-13n+\frac{169}{4}-\frac{25}{4}\right)\\ -250\left(\left(n-\frac{13}{2}\right)^2-\frac{25}{4}\right)\\ -250\left(n-\frac{13}{2}\right)^2+\frac{3125}{2}\]

OpenStudy (bigbosssaint21):

so the last line is the completed square? cause ill need to put it into vertex form

OpenStudy (anonymous):

Yeah, that's the vertex form. \[a(x-h)^2+k\] with vertex \((h,k)\).

OpenStudy (bigbosssaint21):

ok thanks!

OpenStudy (anonymous):

yw

OpenStudy (bigbosssaint21):

um also what is the axis of symmetry of the function P(n) = -250n2 + 3,250n - 9,000

OpenStudy (bigbosssaint21):

if that makes any sense lo

OpenStudy (bigbosssaint21):

*lol

OpenStudy (bigbosssaint21):

@SithsAndGiggles

OpenStudy (bigbosssaint21):

@Jumperman

OpenStudy (bigbosssaint21):

help please

OpenStudy (jumperman):

Okay hold on.

OpenStudy (bigbosssaint21):

sure